OpenStudy (owlcoffee):

I need a hand with the following limit:

OpenStudy (owlcoffee):

\[\lim_{x \rightarrow +\infty} -x ^{2}+\ln \left| x-1 \right|\]

OpenStudy (anonymous):

intuitively it will by \(-\infty\) because \(-x^2\) goes to \(-\infty\) much much faster than \(\log(x-1)\) goes to infinity, but i guess you need a proof of some kind right/

OpenStudy (owlcoffee):

No, thank you, I always have a little trouble with these limits. I completely forgot about the ordenals, thanks Satelite.

OpenStudy (anonymous):

yw

OpenStudy (owlcoffee):

What if I evaluate it when x-> - infinite ?